OpenStudy (anonymous):

Factor the trinomial: (x-4)^2+3(x-4)-18

myininaya (myininaya):

u^2+3u-18 <---do you know how to factor this?

OpenStudy (hesan):

(x-4){(x-4)+3}-18 (x-4){x-1}-18

OpenStudy (anonymous):

@myininaya yes. (u+6)(u-3)

myininaya (myininaya):

great replace u with x-4

OpenStudy (anonymous):

expand your first bracket you get x^2-8x+16 then +3x-!2(this from expanding your second bracket) -18 then you end up with x^2-5x-14 the you use the quadratic formula to find out the value of x.. i ended up with 7 and -2

myininaya (myininaya):

\[(x-4)^2+3(x-4)-18\] Replace x-4 with u \[u^2+3u-18\] (u+6)(u-3) Replace u back with x-4.

OpenStudy (anonymous):

@myininaya so I'd have [(x-4)+6]^2+3[(x-4)-3]-18?

myininaya (myininaya):

(u+6)(u-3) is factored form Replace the u's here with x-4 like this: (x-4+6)(x-4-3) The simplify inside the parenthesis.

myininaya (myininaya):

Then*

OpenStudy (anonymous):

Okay. What I wrote wasn't simplified. So I just needed to factor using substitution. Okay. I started out doing that, but I don't think I was doing it correctly. Thank you much!!

myininaya (myininaya):

Do you want me to give you another one to try? And also don't forget to write -4+6 as 2 and -4-3 as -7 The final answer would be (x+2)(x-7)

myininaya (myininaya):

If you want to, try to factor this one. \[(x-1)^2+5(x-1)+6\]

OpenStudy (anonymous):

Okay, give me a second :)

OpenStudy (anonymous):

u^2+5u+6 (u+3)(u+2) (x-1+3)^2+5(x-1+2)+6 2x^2+5x+6 ... uhhhm I'm stuck. Haha.

myininaya (myininaya):

Now you are doing something weird that you were doing before. Get rid of that third line and everything after.

myininaya (myininaya):

Replace u with x-1 so we will be replacing the u's in (u+3)(u+2) with [x-1] (x-1+3)(x-1+2) = (x+2)(x+1)

OpenStudy (anonymous):

@myininaya you are correct with you (x-7)(x+2)

OpenStudy (anonymous):

Ohh... Oops. Sorry.

myininaya (myininaya):

\[(kitty+bat)^2+3(kitty+bat)+2\] replace kitty+bat with u \[u^2+3u+2=(u+2)(u+1)\] So to factor \[(kitty+bat)^2+3(kitty+bat)+2\] we say \[(kitty+bat+2)(kitty+bat+1)\]

OpenStudy (anonymous):

Ahh. I try to do things to fast. My brain gets ahead of itself... gets all jumbled. Haha, it made sense before kitty+bat, but that's a good way to remember it. Thank you :)

myininaya (myininaya):

lol. I'm just saying whatever those things are as long as they are the same you can always replace with a simple substitution such as u and factor and then go back and replace u with what it was before.

OpenStudy (anonymous):

Yes, I see. XD

myininaya (myininaya):

\[(x^2-1)^2+4(x^2-1)+4\] So we know how to factor: \[u^2+4u+4\]

myininaya (myininaya):

We say \[(u+2)^2\] now do you see what u should be here?

OpenStudy (anonymous):

Humma? You complicated it with x^2 in each of those. Give me a second.

OpenStudy (anonymous):

Ohhhh. Okay.

myininaya (myininaya):

Well instead of x^2-1 we see u so we replaced x^2-1 with u so we can go back and say well u was x^2-1. So ... \[(u+2)^2=(x^2-1+2)^2=(x^2+1)^2\]

OpenStudy (anonymous):

Blahhh. This page keeps freezing on me. I've been trying to type a reply for the last five minutes.

OpenStudy (anonymous):

Yesh. I got that. :/

myininaya (myininaya):

Was that going to be your answer? :)

OpenStudy (anonymous):

Twas...

OpenStudy (anonymous):

I even have it written down. :)

myininaya (myininaya):

Good job. :)

OpenStudy (anonymous):

Thank you for helping me :) I really appreciate it.

OpenStudy (anonymous):

Posting... posting... posting... lol.

OpenStudy (anonymous):

My computer hates this site.

myininaya (myininaya):

I don't think yours is the only one.

myininaya (myininaya):

There is a little lag. I think it might be from so many people using it. I'm not sure.